OpenStudy (anonymous):

Simplify √162 - √160 - 5√32 - √90

OpenStudy (amistre64):

if we could get them to have the same radicals, then its just like working with a like variable

OpenStudy (anonymous):

what is a radical?

OpenStudy (amistre64):

you might be able to get it into 2 term, but I dont see it as one term

OpenStudy (amistre64):

\[\sqrt{\#}\]

OpenStudy (amistre64):

\[\sqrt{160}=\sqrt{16}\sqrt{10}=4\sqrt{10}\] \[\sqrt{90}=\sqrt{9}\sqrt{10}=3\sqrt{10}\] \[-4\sqrt{10}-3\sqrt{10}=(-4-9)\sqrt{10}=-13\sqrt{10}\]

OpenStudy (amistre64):

the other 2 im sure would simplify to a like radical of their own

OpenStudy (anonymous):

I'm sorry but your answer isnt one I can choose from

OpenStudy (amistre64):

thats becasue I only worked out part of it to demonstrate how it would be done, you are spose to take the initiative to complete it or ask questions about anything that may be unclear.